# Synchronization of lasers with complex spatio-temporal dynamics

### v1

@@ Abstract
Synchronization is among the most fascinating dynamical behaviors through which oscillators like lasers align their temporal dynamics under weak coupling. So far, experimental studies have been restricted to lasers exhibiting structured temporal dynamics, such as periodic, mode-locked, or chaotic regimes, the latter usually induced by feedback or optical injection. We now successfully achieve synchronization of two spatially extended lasers with intrinsically chaotic dynamics. We observe synchronization when a dominant transverse mode of the master laser becomes spectrally aligned with a mode of the slave, without requiring matching between the spatial field profiles. Several dynamical regimes are identified, including synchronization of fast chaotic fluctuations and low-frequency polarization hopping, along with switching between normal and inverse synchronization. These results show that the presence of complex dynamics in space and time does not hinder synchronization; instead, it enables a variety of synchronization regimes. Our findings offer new insights into how multimode chaotic lasers can self-organize through weak coupling, highlighting the role of spectral alignment and dominant modes in the emergence of synchronized behavior. In addition, synchronizing lasers that are chaotic without feedback enables simpler and more compact systems in view of applications such as physical cryptography or optical computing.
@@ I Introduction
More generally, synchronization means the ability of weakly coupled systems to develop correlated temporal dynamics. Synchronization can also occur between coupled chaotic systems, in spite of their inherent sensitivity to initial conditions [10]. Semiconductor lasers are an outstanding platform to study chaos synchronization due to their rich dynamical behavior under optical injection and feedback [11, 12, 13, 14, 15, 16]. Over the past decades, numerous forms of synchronization, such as complete, phase, frequency, generalized, and even anticipatory synchronization, have been explored [17, 18, 19, 20, 21]. Investigations into the coupling mechanisms have revealed the critical roles of both local and global interactions in shaping the stability and diversity of synchronization regimes, leading to applications like secure communication based on chaos encryption [17, 22, 23] and photonic computing schemes exploiting collective dynamics for information processing [24].
While temporal synchronization has been thoroughly explored, the extension to spatio-temporal systems remains largely unexplored. Systems with significant transverse extent can exhibit complex space- and time-dependent dynamics, and the spatial dimension introduces new challenges for achieving synchronization. Unlike purely temporal oscillators, spatially extended systems are prone to pattern formation, filamentation [25, 26], and modal competition [27], which can destabilize synchronization. Theoretical studies have addressed synchronization in multimode and distributed optical systems [28, 29], but despite its potential for parallel chaotic information processing and multichannel transmission [30, 31], spatio-temporal chaos synchronization has seen limited experimental validation. So far, synchronization of complex spatio-temporal dynamics has been experimentally demonstrated only using a liquid crystal light valve with feedback [32].
In order to address these fundamental questions experimentally, we take advantage of the intrinsically complex spatio-temporal dynamics of broad-area vertical-cavity surface-emitting lasers (BA-VCSELs) [33, 34, 35, 36, 37]. The mode competition in these devices naturally results in sustained high-frequency fluctuations and chaos without requiring external perturbations like feedback [38, 39]. The main question is if chaotic broad-area lasers can achieve synchronization through weak optical coupling.
@@
Figure 1: (a) Schematic representation of the synchronization experiment setup. This configuration allows both the analysis of the dynamics of the master laser, and the response of the slave laser under injection. It is possible to study both of the lasers in the spatio-spectral domain, using an imaging spectrometer (with a half-wave plate (HWP) to control the polarization orientation at the spectrometer input), or in the temporal domain, using two high-bandwidth photodetectors connected to a high-bandwidth real-time oscilloscope. (b1) Spatio-spectral images of the free-running master laser for different pump currents in $u$-polarization, which is selected for injection; (b2) Spatio-spectral image in both $u$- and $v$-polarizations for $I=6.5$ mA, with the strong modes $M_{u,v}(3,1)$ highlighted. The symmetry axis of the transverse mode is given by the blue dashed line to indicate the + (-) orientation of the modes. Master laser dynamics for $I_{M}$= 6 mA (c1) and $I_{M}$= 9 mA (c2) in the time and frequency domains. The vertical dashed lines in the RF-spectra indicate the birefringence $\Delta\nu_{b}$. The red time trace in (c1) is low-pass filtered at $0.1$ GHz.
@@
Specifically, we investigate two coupled BA-VCSELs which exhibit chaotic dynamics in free-running operation (see Fig. 1). Via the pump current, we can change the number of transverse lasing modes and the dynamical regimes [38, 39]. Synchronization of chaotic dynamics is demonstrated for weak coupling via unidirectional optical injection. Remarkably, synchronization emerges despite significant differences in the spatial mode structures. It is mediated by spectral alignment between strong transverse modes, establishing a mechanism where frequency alignment prevails over spatial matching in enabling chaos synchronization in multimode systems.
Our results open new perspectives for spatially extended optical systems with chaotic dynamics, with implications for high-capacity secure communications and massively-parallel chaotic information processing. They establish that spatial complexity, instead of preventing synchronization, can coexist with robust collective dynamics under appropriate coupling conditions.
The paper is organized as follows. We first describe the experimental setup used for unidirectional coupling between two BA-VCSELs. We then characterize the spontaneous chaotic dynamics of these lasers in free-running operation. Synchronization results are subsequently presented for two distinct regimes: a first regime involving slow polarization-hopping dynamics, and a second regime featuring broadband chaotic dynamics. We conclude by discussing the implications of these results and the perspectives they open for the study of complex spatio-temporal systems.
II Results
@@
We experimentally study unidirectional optical injection of one BA-VCSEL into another BA-VCSEL of the same model and analyze the dynamical response of the slave laser to the signal of the master laser. A simplified schematic of the experimental setup is shown in Fig. 1(a), and a detailed description is provided in Appendix 1. We begin with a summary of the properties of the master laser (see also Refs. [38, 39]), and a comparison to the slave laser is presented in Appendix 3.
@@
The free-running BA-VCSELs lase in several transverse modes and two linear orthogonal polarization states, denoted by $u$ and $v$. The dominant polarization depends on the pump current and changes at several polarization-switching points (PSPs, see Appendix Fig. A5) [38]. The birefringence splitting between $u$- and $v$- polarization modes is around $\Delta\nu_{b}\approx 9$ GHz (see Appendix 3), and we refer to ”red” ($u$) and ”blue” ($v$) polarization according to their relative position in the optical spectrum.
@@
Figure 1(b) shows spatio-spectral images measured with an imaging spectrometer (see Appendix 1) for different pump currents. With increasing pump current, more transverse modes are progressively excited. A transverse mode with mode indices $(m,n)$ has $2m$ intensity maxima in azimuthal direction and $n$ maxima in radial direction. Furthermore, for $m>0$, two distinct spatial orientations are possible, which we denote with $+$ (--) when they are (anti-) symmetric with respect to a symmetry axis as shown in Fig. 1(b2). In the following we denote the modes with $X(m,n,\pm)$ where $X\in\{M,S\}$ refers to the master and slave laser, respectively. It should be noted that modes $(m,n,+)$ and $(m,n,-)$ are not necessarily degenerate [40], but can exhibit a splitting $\Delta\nu_{O}$ of the order of several GHz (see Appendix A.4). The actual splitting seems to depend on the transverse mode. Typically, only one of the orientations is excited for a given polarization, and the corresponding mode in the other polarization has the opposite orientation due to gain competition [41, 42, 43, 38]. This is exemplified in Fig. 1(b2) with the modes $M_{v}(3,1,+)$ and $M_{u}(3,1,-)$. However, in some cases both orientations of a transverse mode $(m,n)$ lase in the same polarizations as shown in Appendix Fig. A8.
@@
The competition of lasing modes with different spatial profiles and polarizations results in rich nonlinear dynamics, as shown in Fig. 1(c). As the pump current increases, the system undergoes a sequence of bifurcations that progressively enrich the temporal behavior [38, 39]: initially periodic oscillations transition to quasi-periodic regimes, and eventually to chaotic dynamics (see Appendix Fig. A9). These bifurcations are usually accompanied by a redistribution of power between different transverse modes and in some cases polarization switching points [38]. We restrict our analysis to parameter regions in which the master laser features chaotic dynamics for a focused investigation of synchronization in the presence of intrinsic multimode chaos.
@@
The lasing dynamics features different time scales: the dominant frequencies are around the birefringence $\Delta\nu_{b}$ [see Fig. 1(c)]. However, in some current regimes like around $6$ mA, this fast dynamics coexists with a slower polarization-hopping dynamics with frequencies of the order of $100$ MHz [see Fig. 1(c1) and Appendix Fig. A7], similar to observations for single-mode VCSELs [44, 45, 46, 47].
@@
Next, we implement optical injection between two BA-VCSELs to investigate synchronization between them. Although the two lasers are of the same model, some differences in their spectral and dynamical properties inevitably remain. The comparison of their optical spectra and dynamics in Appendix B shows that spectral alignment can be achieved by adjusting the detuning via the temperature or the pump current of the slave (see Appendix C). The two distinct dynamical regimes in Fig. 1(c) are injected into the slave laser in order to analyze the synchronization quality for both high frequency chaos and slower mode-hopping dynamics. Different synchronization mechanisms emerge depending on the properties of the injected dynamics. The response of the slave laser strongly depends on the temporal and spectral structure of the injected signal, highlighting the richness of coupling phenomena in chaotic multimode systems.
@@
Figure 2: Spatio-spectral images in $u$-polarization of master and slave laser for $\Delta\nu=$ 0 GHz (a1), 80 GHz (a2), and 133 GHz (a3). (b1) Correlation between the unfiltered and (b2) low-pass filtered (cutoff at 1 GHz) time traces in $u-$polarization of master and slave laser as a function of the detuning or pump current $I_{S}$. (c) Low-pass filtered time traces (cutoff at 0.1 GHz) of the master (black) and slave (red) lasers for $\Delta\nu=0$ GHz (top) and $\Delta\nu=1.7$ GHz (bottom), corresponding to a correlation of approximately $\pm$80%, respectively. (d) Absolute value of correlation for $\Delta\nu=0$ GHz (blue), $\Delta\nu=1.7$ GHz (green), and $\Delta\nu=180$GHz (red) as a function of the low-pass filter cutoff frequency. The vertical blue dashed lines indicate the cutoff value chosen to compute the correlation coefficient shown in (b2).
@@
The first case we study is when the master laser operates at $I_{M}=6.17$ mA and exhibits a dynamics characterized by both high-frequency components and slow polarization-hopping (typically around $100$ MHz). The master laser current and temperature are fixed to keep its dynamical properties constant. The slave laser current is varied between $I_{S}=2$ mA and $10$ mA at a fixed temperature of $20^{\circ}$. The slave laser modes are red-shifted due to Joule heating as its current increases [see Appendix C]. Furthermore, its output power increases, thereby reducing the optical injection ratio (see Appendix A). We study parallel injection, meaning that the $u$-polarization component of the master laser, which is its dominant polarization for this pump current, is injected into the $u$-polarization of the slave laser using a HWP. The slave laser has several PSPs in this current range, but its red polarization ($u$) remains mostly dominant (see Appendix Fig. A5).
@@
Figure 2(b1) shows the correlation between the time traces (see Appendix A) of $u$-polarized emissions of master and slave laser as function of the frequency detuning $\Delta\nu$ between the two lasers, which is defined here as the frequency of mode $M_{u}(0,1)$ minus the frequency of $S_{u}(0,1)$ as illustrated in Fig. 2(a1). We observe several cases of significant positive and negative values of correlation, corresponding to synchronization and inverse synchronization respectively, which are discussed in the following.
@@
The positive correlation peaks are related to spectral alignment between transverse modes of the master and the slave laser with high power. For $\Delta\nu=0$, the spatio-spectral image in Fig. 2(a1) shows that both the $S_{u}(0,1)$ and the $S_{u}(3,1)$ modes emit strongly and are spectrally aligned with the $M_{u}(0,1)$ and $M_{u}(3,1)$ modes, respectively. We originally expected this to yield the best synchronization since theoretically each transverse mode of the master laser would be aligned with its counterpart of the slave, but in practice perfect matching of all transverse mode frequencies is not achievable due to small differences in modal spacings between the two lasers. We also observe that for $\Delta\nu=0$, when the $S_{u}(3,1)$ mode aligns with the $M_{u}(3,1)$ mode, its power significantly increases while neighboring modes become weaker, as discussed in Appendix Fig. A13. This highlights the strong interaction of these two modes as their frequencies align.
@@
Similar observations are made for the other positive correlation peaks. At $\Delta\nu\approx 30$ GHz, a weak synchronization peak appears, associated with the alignment of $M_{u}(0,1)$ with $S_{u}(1,1)$. The peak at $\Delta\nu\approx 80$ GHz corresponds to the alignment between $M_{u}(0,1)$ and $S_{u}(2,1)$ [see Fig. 2(a2) and Appendix Fig. A13]. Similarly, at $\Delta\nu\approx 133$ GHz, another peak emerges from the alignment between modes $M_{u}(0,1)$ and $S_{u}(3,1)$ [see Fig. 2(a2)]. These examples demonstrate that synchronization can be achieved when a strong transverse mode of the master laser spectrally aligns with a transverse mode of the slave, however, it need not be the same transverse modes: spectral alignment appears to be more important than spatial alignment for successful synchronization. It should also be emphasized that the other transverse modes of the slave laser remain active and are different from those of the master laser [see Fig. 2(a)], meaning that synchronizing the temporal dynamics of the lasers does not require synchronizing the optical spectrum or the spatial intensity distribution.
@@
The first example at $\Delta\nu=1.7$ GHz happens right after the alignment of the $M_{u}(3,1)$ mode with the $S_{u}(3,1)$ mode when the former approaches the frequency of the $S_{v}(3,1)$ mode, as shown by the time traces of the master and slave, low-pass filtered at 0.1 GHz cutoff, in Fig. 2(c). Indeed, Appendix Fig. A11(a) shows that the $S_{v}(3,1)$ mode is strongly excited at $\Delta\nu=1.7$ GHz. We surmise that when injecting light into the $u$-polarization at a frequency near a $v$-polarized mode of the slave, it can create a synchronization of the $v$-polarized emission with the mode-hopping dynamics of the $u$-polarized injection signal from the master, though this is evidently not always the case. Since the emission in $u$- and $v$-polarization is highly anticorrelated in the polarization-hopping regime [see Appendix Fig. A7], the $u$-polarized emission of the slave becomes anti-correlated to the $u$-polarized master signal since the latter is synchronized with the $v$-polarized emission of the slave. At $180$ GHz detuning, we observe that the $S_{v}(1,2)$ mode is enhanced when its frequency comes close to the $M_{v}(0,1)$ mode.
@@
These examples demonstrate that coupling of two multimode VCSELs can create different types of synchronization depending on system parameters, and that moreover strong dynamic correlations can be established even for detunings as large as $180$ GHz.
@@
Finally we analyze the master-slave correlation across different timescales by applying spectral filtering (see Appendix B). Figure 2(d) shows the evolution of the correlation as function of the low-pass filter cutoff frequency for three examples. The correlation significantly improves with correlations of up to 90% for $100$ MHz cutoff at $\Delta\nu$ = 180 GHz. Furthermore, Fig. 2(b2) shows the correlation of the low-pass filtered time traces with 1 GHz cutoff as a function of detuning. A global increase is observed across all correlation regions. These low-frequency components represent the relatively slow polarization-hopping dynamics which appears to synchronize much better than the high-frequency components of the dynamics as demonstrated by the low-pass filtered time traces of master and slave laser at $\Delta\nu=0$ and $1.7$ GHz in Fig. 2(c). This demonstrates that it is the polarization-hopping dynamics which is synchronized, and that one can achieve very high synchronization quality even though the BA-VCSELs are not perfectly identical.
@@
Figure 3: (a,b) Spatio-spectral images of $u$-polarized emission of the slave laser under injection (top) and the master laser (bottom) for $I_{M}=8.8\ \mathrm{mA}$ and $I_{S}=4.8\ \mathrm{mA}$. The detunings are $\Delta\nu=$3.50 GHz (a1), 0 GHz (a2), 4.14 GHz (a3), 40 GHz (b1), 44.3 GHz (b2), and 48.0 GHz (b3). The symmetry axis of the transverse modes is indicated with the blue dashed lines. (c) Correlation between the $u-$polarized time traces of master and slave laser as a function of the detuning, without filter (green) and with low-pass filter at 1 GHz cutoff (black). The magnification on the right shows the three-peak scenario around $\Delta\nu=$ 44.3 GHz.
@@
For the second case of injection that we study, the master laser is set to $I_{M}=8.8$ mA and $T_{M}=20^{\circ}$, the slave laser is operated at a fixed current of $I_{S}=4.8$ mA, and the detuning is varied via the temperature of the slave. As before, the master laser emits predominantly along the $u$-polarization, and parallel injection of the $u$-polarization is performed. These experimental parameters lead to three differences to the first case. First, the injection ratio is higher though it remains relatively weak (see Appendix A). Second, while the master laser has higher power, it is distributed over a larger number of transverse modes [see Fig. 3(a)]. Third, the master laser operates in a chaotic state with high frequency components (see Fig. 1(c) and Ref. [39]) without polarization-hopping. In the following we discuss how these changes affect the synchronization.

### v2

# arXiv:2506.15351v2[physics.optics] 02 Mar 2026

@@ Abstract
Synchronization is the spontaneous alignment of the dynamics of weakly-coupled oscillators. In addition to temporal dynamics like periodic and chaotic oscillations, also the spatio-temporal dynamics of spatially-extended systems like wildlife populations can synchronize. We exploit here the intrinsic spatio-temporal complex dynamics of broad area lasers to demonstrate such synchronization at lab-scale. Broad-area vertical-cavity surface-emitting lasers (BA-VCSELs) exhibit chaos from the nonlinear coupling between laser modes with different spatial profiles and polarization. When coupling two BA-VCSELs, several synchronization and anti-synchronization regimes are observed, highlighting the complex interplay between oscillating modes with different frequencies and spatial patterns. The correlation coefficient varies between 0.2 and 0.9 depending on the dynamics and on the time scale under analysis. Besides its fundamental interest, our experiment with commercial devices marks the first step towards real-world spatial multiplexing in multiple user physical-layer secure communication based on chaos synchronization.
@@ Introduction
More generally, synchronization means the ability of weakly coupled systems to develop correlated temporal dynamics. Synchronization can also occur between coupled chaotic systems, in spite of their inherent sensitivity to initial conditions [41, 43, 46, 5]. Chaos synchronization is typically achieved with the master-slave unidirectional coupling configuration [36, 42], but also common-signal induced synchronization has been reported. [52, 48]. While temporal synchronization has been thoroughly explored, the extension to spatio-temporal systems remains largely unexplored. Systems with significant transverse extent can exhibit complex space- and time-dependent dynamics, but the spatial dimension introduces new challenges for achieving and understanding synchronization.
Synchronization of complex spatio-temporal dynamics occurs naturally for example in ecological systems [13, 12], the human brain [7] and weather patterns [34]. Understanding and replicating these dynamics in artificial systems and at lab scale could revolutionize neuroscience, machine learning, biology and our understanding of complex systems in general.
Optical systems are an ideal testbed for studying nonlinear dynamics thanks to their compact size and fast time scales [43]. Complex spatio-temporal dynamics like pattern formation, filamentation [1, 20], and modal competition [9] can be found for example in broad-area semiconductor lasers. In addition, theoretical studies have predicted the possibility to synchronize multimode and distributed optical systems [17, 22] and shown their potential for parallel chaotic information processing and multiple user encryption schemes [50, 6]. Nonetheless synchronization of complex spatio-temporal dynamics has been experimentally demonstrated so far only using a liquid crystal light valve with feedback [24], and its implementation therefore remains far from practical physical-layer communication architectures.
We investigate the possibility to synchronize spatio-temporal chaotic systems using commercially available lasers in a table-top experiment. We take advantage of our recent discovery [11, 35] that broad area vertical cavity surface-emitting lasers (BA-VCSELs) naturally exhibit chaos involving nonlinear interactions between a large number of spatial modes with different polarization states [45, 8, 23, 15, 16]. By coupling two such lasers, we demonstrate synchronization in a controllable and reproducible way and analyze how synchronization relates to the spatial lasing modes. Besides its fundamental interest for the study of complex systems, our work marks the first step to a realistic implementation of spatial multiplexing for physical-layer secure communications at high bit rates.
@@
Figure 1: Setup and free-running master laser. (a) Schematic representation of the synchronization experiment setup. This configuration allows both the analysis of the dynamics of the master laser, and the response of the slave laser under injection. It is possible to study both of the lasers in the spatio-spectral domain, using an imaging spectrometer (with a half-wave plate (HWP) to control the polarization orientation at the spectrometer input), or in the temporal domain, using two high-bandwidth photodetectors connected to a high-bandwidth real-time oscilloscope. (b1) Spatio-spectral images of the free-running master laser for different pump currents in $u$-polarization, which is selected for injection; (b2) Spatio-spectral image in both $u$- and $v$-polarizations for $I=6.5$ mA, with the strong modes $M_{u,v}(3,1)$ highlighted. The symmetry axis of the transverse mode is given by the blue dashed line to indicate the + (-) orientation of the modes. Master laser dynamics for $I_{M}$= 6 mA (c1) and $I_{M}$= 9 mA (c2) in the time and frequency domains. The vertical dashed lines in the RF-spectra indicate the birefringence $\Delta\nu_{b}$. The red time trace in (c1) is low-pass filtered at $0.1$ GHz.
@@
Specifically, we investigate two coupled BA-VCSELs which exhibit chaotic dynamics in free-running operation (see Fig. 1). Via the pump current, we can change the number of transverse lasing modes and the dynamical regimes [11, 35]. Synchronization of chaotic dynamics is demonstrated for weak coupling via unidirectional optical injection. Remarkably, synchronization emerges despite significant differences in the spatial mode structures. It is mediated by the spectral alignment of the strong transverse modes of the master laser, defined here as those containing a significant portion of its total power. This reveals a mechanism in which frequency alignment prevails over spatial matching in enabling chaos synchronization in multimode systems. Spatial complexity, instead of preventing synchronization, therefore coexists with robust collective dynamics under appropriate coupling conditions.
We propose to investigate the effect of unidirectional optical injection between two BA-VCSELs in two distinct dynamical regimes: a first regime involving slow polarization-hopping dynamics, and a second regime featuring broadband chaotic dynamics.
Results
@@
We experimentally study unidirectional optical injection of one BA-VCSEL into another BA-VCSEL of the same model and analyze the dynamical response of the slave laser to the signal of the master laser. A simplified schematic of the experimental setup is shown in Fig. 1(a), and a detailed description is provided in Supplementary Section 1. We begin with a summary of the properties of the master laser (see also Refs. [11, 35]), and a comparison to the slave laser is presented in Supplementary Section 2.
@@
The free-running BA-VCSELs lase in several transverse modes and two linear orthogonal polarization states, denoted by $u$ and $v$. The dominant polarization depends on the pump current and changes at several polarization-switching points (PSPs, see Supplementary Fig. S2) [11]. The birefringence splitting between $u$- and $v$- polarization modes is around $\Delta\nu_{b}\approx 9$ GHz (see Supplementary Section 2), and we refer to ”red” ($u$) and ”blue” ($v$) polarization according to their relative position in the optical spectrum.
@@
Figure 1(b) shows spatio-spectral images measured with an imaging spectrometer (see Supplementary Section 1) for different pump currents. With increasing pump current, more transverse modes are progressively excited. A transverse mode with mode indices $(m,n)$ has $2m$ intensity maxima in azimuthal direction and $n$ maxima in radial direction. Furthermore, for $m>0$, two distinct spatial orientations are possible, which we denote with ++ (-) when they are (anti-) symmetric with respect to a symmetry axis as shown in Fig. 1(b2). In the following we denote the modes with $X(m,n,\pm)$ where $X\in\{M,S\}$ refers to the master and slave laser, respectively. It should be noted that modes $(m,n,+)$ and (m,n,)(m,n,-) are not necessarily degenerate [30], but can exhibit a splitting $\Delta\nu_{O}$ of the order of several GHz (see Supplementary Section A.4). The actual splitting seems to depend on the transverse mode. Typically, only one of the orientations is excited for a given polarization, and the corresponding mode in the other polarization has the opposite orientation due to gain competition [19, 31, 33, 11]. This is exemplified in Fig. 1(b2) with the modes $M_{v}(3,1,+)$ and $M_{u}(3,1,-)$. However, in some cases both orientations of a transverse mode $(m,n)$ lase in the same polarizations as shown in Supplementary Fig. S5.
@@
The competition of lasing modes with different spatial profiles and polarizations results in rich nonlinear dynamics, as shown in Fig. 1(c). As the pump current increases, the system undergoes a sequence of bifurcations that progressively enrich the temporal behavior [11, 35]: initially periodic oscillations transition to quasi-periodic regimes, and eventually to chaotic dynamics (see Supplementary Fig. S6). These bifurcations are usually accompanied by a redistribution of power between different transverse modes and in some cases polarization switching points [11]. We restrict our analysis to parameter regions in which the master laser features chaotic dynamics for a focused investigation of synchronization in the presence of intrinsic multimode chaos.
@@
The lasing dynamics features different time scales: the dominant frequencies are around the birefringence $\Delta\nu_{b}$ [see Fig. 1(c)]. However, in some current regimes like around $6$ mA, this fast dynamics coexists with a slower polarization-hopping dynamics with frequencies of the order of $100$ MHz [see Fig. 1(c1) and Supplementary Fig. S4], similar to observations for single-mode VCSELs [32, 51, 47, 39].
@@
Next, we implement optical injection between two BA-VCSELs to investigate synchronization between them. Although the two lasers are of the same model, some differences in their spectral and dynamical properties inevitably remain. The comparison of their optical spectra and dynamics in Supplementary Section B shows that spectral alignment can be achieved by adjusting the detuning via the temperature or the pump current of the slave (see Supplementary Section C). The two distinct dynamical regimes in Fig. 1(c) are injected into the slave laser in order to analyze the synchronization quality for both high frequency chaos and slower mode-hopping dynamics. Different synchronization mechanisms emerge depending on the properties of the injected dynamics. The response of the slave laser strongly depends on the temporal and spectral structure of the injected signal, highlighting the richness of coupling phenomena in chaotic multimode systems.
@@
Figure 2: Synchronization of polarization-hopping dynamics. Spatio-spectral images in $u$-polarization of master and slave laser for $\Delta\nu=$ 0 GHz (a1), 80 GHz (a2), and 133 GHz (a3). The camera field of view corresponds to 2600 $\mu m$ $\times$ 740 $\mu m$ in the object plane. (b1) Correlation between the unfiltered and (b2) low-pass filtered (cutoff at 1 GHz) time traces in $u$-polarization of master and slave laser as a function of the detuning or pump current $I_{S}$. (c) Low-pass filtered time traces (cutoff at 0.1 GHz) of the master (black) and slave (red) lasers for $\Delta\nu=0$ GHz (top) and $\Delta\nu=1.7$ GHz (bottom), corresponding to a correlation of approximately $\pm$80%, respectively. (d) Absolute value of correlation for $\Delta\nu=0$ GHz (blue), $\Delta\nu=1.7$ GHz (green), and $\Delta\nu=180$ GHz (red) as a function of the low-pass filter cutoff frequency. The vertical blue dashed lines indicate the cutoff value chosen to compute the correlation coefficient shown in (b2).
@@
The first case we study is when the master laser operates at $I_{M}=6.17$ mA and exhibits a dynamics characterized by both high-frequency components and slow polarization-hopping (typically around $100$ MHz). The master laser current and temperature are fixed to keep its dynamical properties constant. The slave laser current is varied between $I_{S}=2$ mA and $10$ mA at a fixed temperature of $20^{\circ}$C. The slave laser modes are red-shifted due to Joule heating as its current increases [see Supplementary Section C]. Furthermore, its output power increases, thereby reducing the optical injection ratio (see Supplementary Section A). We study parallel injection, meaning that the $u$-polarization component of the master laser, which is its dominant polarization for this pump current, is injected into the $u$-polarization of the slave laser using a HWP. The slave laser has several PSPs in this current range, but its red polarization ($u$) remains mostly dominant (see Supplementary Fig. S2).
@@
Figure 2(b1) shows the correlation between the time traces (see Methods) of $u$-polarized emissions of master and slave laser as function of the frequency detuning $\Delta\nu$ between the two lasers, which is defined here as the frequency of mode $M_{u}(0,1)$ minus the frequency of $S_{u}(0,1)$ as illustrated in Fig. 2(a1). We observe several cases of significant positive and negative values of correlation, corresponding to synchronization and inverse synchronization respectively, which are discussed in the following.
@@
The positive correlation peaks are related to spectral alignment between transverse modes of the master and the slave laser with high power. For $\Delta\nu=0$, the spatio-spectral image in Fig. 2(a1) shows that both the $S_{u}(0,1)$ and the $S_{u}(3,1)$ modes emit strongly and are spectrally aligned with the $M_{u}(0,1)$ and $M_{u}(3,1)$ modes, respectively. We originally expected this to yield the best synchronization since theoretically each transverse mode of the master laser would be aligned with its counterpart of the slave, but in practice perfect matching of all transverse mode frequencies is not achievable due to small differences in modal spacings between the two lasers. We also observe that for $\Delta\nu=0$, when the $S_{u}(3,1)$ mode aligns with the $M_{u}(3,1)$ mode, its power significantly increases while neighboring modes become weaker, as discussed in Supplementary Fig. S10. This highlights the strong interaction of these two modes as their frequencies align.
@@
Similar observations are made for the other positive correlation peaks. At $\Delta\nu\approx 30$ GHz, a weak synchronization peak appears, associated with the alignment of $M_{u}(0,1)$ with $S_{u}(1,1)$. The peak at $\Delta\nu\approx 80$ GHz corresponds to the alignment between $M_{u}(0,1)$ and $S_{u}(2,1)$ [see Fig. 2(a2) and Supplementary Fig. S10]. Similarly, at $\Delta\nu\approx 133$ GHz, another peak emerges from the alignment between modes $M_{u}(0,1)$ and $S_{u}(3,1)$ [see Fig. 2(a2)]. These examples demonstrate that synchronization can be achieved when a strong transverse mode of the master laser spectrally aligns with a transverse mode of the slave, however, it need not be the same transverse modes: spectral alignment appears to be more important than spatial alignment for successful synchronization. It should also be emphasized that the other transverse modes of the slave laser remain active and are different from those of the master laser [see Fig. 2(a)], meaning that synchronizing the temporal dynamics of the lasers does not require synchronizing the optical spectrum or the spatial intensity distribution.
@@
The first example at $\Delta\nu=1.7$ GHz happens right after the alignment of the $M_{u}(3,1)$ mode with the $S_{u}(3,1)$ mode when the former approaches the frequency of the $S_{v}(3,1)$ mode, as shown by the time traces of the master and slave, low-pass filtered at 0.1 GHz cutoff, in Fig. 2(c). Indeed, Supplementary Fig. S8(a) shows that the $S_{v}(3,1)$ mode is strongly excited at $\Delta\nu=1.7$ GHz. We surmise that when injecting light into the $u$-polarization at a frequency near a $v$-polarized mode of the slave, it can create a synchronization of the $v$-polarized emission with the mode-hopping dynamics of the $u$-polarized injection signal from the master, though this is evidently not always the case. Since the emission in $u$- and $v$-polarization is highly anticorrelated in the polarization-hopping regime [see Supplementary Fig. S4], the $u$-polarized emission of the slave becomes anti-correlated to the $u$-polarized master signal since the latter is synchronized with the $v$-polarized emission of the slave. At $180$ GHz detuning, we observe that the $S_{v}(1,2)$ mode is enhanced when its frequency comes close to the $M_{v}(0,1)$ mode.
@@
These examples demonstrate that coupling of two multimode VCSELs can create different types of synchronization depending on system parameters, and that moreover strong dynamic correlations can be established between transverse modes of very different order.
@@
Finally we analyze the master-slave correlation across different timescales by applying spectral filtering (see Methods). Figure 2(d) shows the evolution of the correlation as function of the low-pass filter cutoff frequency for three examples. The correlation significantly improves with correlations of up to 90% for $100$ MHz cutoff at $\Delta\nu$ = 180 GHz. Furthermore, Fig. 2(b2) shows the correlation of the low-pass filtered time traces with 1 GHz cutoff as a function of detuning. A global increase is observed across all correlation regions. These low-frequency components represent the relatively slow polarization-hopping dynamics which appears to synchronize much better than the high-frequency components of the dynamics as demonstrated by the low-pass filtered time traces of master and slave laser at $\Delta\nu=0$ and $1.7$ GHz in Fig. 2(c). This demonstrates that it is the polarization-hopping dynamics which is synchronized, and that one can achieve very high synchronization quality even though the BA-VCSELs are not perfectly identical.
@@
Figure 3: Synchronization of broadband chaotic dynamics. (a,b) Spatio-spectral images of $u$-polarized emission of the slave laser under injection (top) and the master laser (bottom) for $I_{M}=8.8\ \mathrm{mA}$ and $I_{S}=4.8\ \mathrm{mA}$. The camera field of view corresponds to 2470 $\mu m$ $\times$ 950 $\mu m$ in the object plan. The detunings are $\Delta\nu=$ -3.50 GHz (a1), 0 GHz (a2), 4.14 GHz (a3), 40 GHz (b1), 44.3 GHz (b2), and 48.0 GHz (b3). The symmetry axis of the transverse modes is indicated with the blue dashed lines. (c) Correlation between the $u$-polarized time traces of master and slave laser as a function of the detuning, without filter (green) and with low-pass filter at 1 GHz cutoff (black). The magnification on the right shows the three-peak scenario around $\Delta\nu=$ 44.3 GHz.
@@
For the second case of injection that we study, the master laser is set to $I_{M}=8.8$ mA and $T_{M}=20^{\circ}$C, the slave laser is operated at a fixed current of $I_{S}=4.8$ mA, and the detuning is varied via the temperature of the slave. As before, the master laser emits predominantly along the $u$-polarization, and parallel injection of the $u$-polarization is performed. These experimental parameters lead to three differences to the first case. First, the injection ratio is higher though it remains relatively weak (see Supplementary Section A). Second, while the master laser has higher power, it is distributed over a larger number of transverse modes [see Fig. 3(a)]. Third, the master laser operates in a chaotic state with high frequency components (see Fig. 1(c) and Ref. [35]) without polarization-hopping. In the following we discuss how these changes affect the synchronization.